OpenStudy (anonymous):

Could someone please help me with this? Variable y varies directly with x2, and y = 96 when x = 4. What is the value of y when x = 2? y = ?

OpenStudy (anonymous):

y = kx^2 Plug in what we know: 96 = 4^2k 96 = 16k Divide 16 to both sides: k = 6 So the constant is 6. Now we can plug this into the equation along with x = 2. y = (6)(2^2) y = (6)(4) Can you multiply 6 * ?

OpenStudy (anonymous):

Can you multiply 6 * 4?*

OpenStudy (anonymous):

Yes I can, thank you for your help!! :D

OpenStudy (anonymous):

No problem!